The length of a mercury column of a thermometer are 1.06cm &20.86cm respectively at a standard fixed points. What is the temperature of the body which produces 7.0cm of this mercury column
1
Expert's answer
2019-09-16T09:40:53-0400

Solution:


Standard fixed points are:


"t_0=0^{\circ} С"


"t_1=100^{\circ} С"


Let's write the linear expansion equation for temperature t1:


"l_1-l_0=\alpha{l_0}(t_1-t_0)"


From previous equation we can find the coefficient of thermal expansion of mercury:


"\alpha=\frac{l_1-l_0}{l_0(t_1-t_0)}=\frac{20.86-1.06}{1.06\cdot{100}}=0.1868" K-1


Let's write the linear expansion equation for temperature t2:


"l_2-l_0=\alpha{l_0(t_2-t_0)}"


"t_2-t_0=\frac{l_2-l_0}{\alpha{l_0}}=\frac{7.0-1.06}{0.1868\cdot{1.06}}=30^{\circ} С"


"t_2=30^{\circ} С"


Answer:

The temperature of the body is 30o C.